A line passes through the points and (-10,-4) (-1,2) What is the slope of the line that passes through the two points?
Answer by ganta.santoshgk@gmail.com(1)

slope of line for (x1,y1), (x2,y2) is
m=(y2-y1)/(x2-x1)
here m=6/(-9)
=-2/3